# Gluon Polarimetry with Energy-Energy Correlators

### v1

@@ Abstract
Introduction— Understanding the linear polarization of gluons inside unpolarized nucleons is essential for mapping their gluonic structure, a topic that has drawn significant interest from both small-$x$ and high-energy spin physics communities Mulders and Rodrigues (2001); Metz and Zhou (2011); Schafer and Zhou (2012); Boer et al. (2009, 2011); Qiu et al. (2011); Pisano et al. (2013); Boer et al. (2016, 2017); Kishore and Mukherjee (2019); Dominguez et al. (2012); Dumitru et al. (2015); Sun et al. (2011); Boer et al. (2012); Catani et al. (2014); Boer and den Dunnen (2014); Echevarria et al. (2015); Ma et al. (2013a); Gutierrez-Reyes et al. (2019); den Dunnen et al. (2014); Zhang (2014); Ma and Wang (2016); Altinoluk et al. (2024); Taels et al. (2022). The recent finding Bhattacharya et al. (2024), together with the observation that gluons exhibit 100% linear polarization at small $x$ in the dilute limit Metz and Zhou (2011), indicate that the helicity and orbital angular momentum (OAM) of gluons are maximally quantum entangled. This phenomenon has broad phenomenological implications, as evidenced by predicted substantial $\cos 2\phi$ azimuthal asymmetries in two particle production processes Boer et al. (2009, 2011); Qiu et al. (2011); Pisano et al. (2013); Boer et al. (2016); Schafer and Zhou (2012); Boer et al. (2017) and the transverse momentum distribution of Higgs boson at LHC Sun et al. (2011); Boer et al. (2012); Catani et al. (2014); Boer and den Dunnen (2014); Echevarria et al. (2015); Ma et al. (2013a); den Dunnen et al. (2014); Zhang (2014); Ma and Wang (2016); Gutierrez-Reyes et al. (2019). Despite these compelling signatures, directly measuring linear gluon polarization remains a major challenge. Traditional methods, such as those based on Transverse Momentum Dependent (TMD) factorization, suffer from two key limitations: the polarization signal is diluted by TMD evolution effects at low transverse momentum, and identical $\cos 2\phi$ asymmetries can also arise from final-state soft gluon emissions contaminating the extraction of Hatta et al. (2021a, b). To address these challenges, emerging observables like Energy-Energy Correlations (EECs) offer a novel and potentially more precise method for measuring linear gluon polarization.
@@
Figure 1: The figure depicts a proton-proton collision in which a nearly collinear, linearly polarized gluon is emitted from an incoming parton involved in a hard scattering event. The magnified view on the right shows the resulting EEC within the jet cone initiated by the radiated gluon, clearly displaying the $\cos 2\phi$ anisotropy as revealed by our numerical calculations.
@@
EEC based methods also leverage the $\cos 2\phi$ azimuthal asymmetry to probe linear gluon polarization. Specifically, a $\cos 2\phi$ azimuthal modulation in the squeezed limit of the three-point energy correlator within a jet Chen et al. (2021, 2022) provides a direct probe of the linear polarization of intermediate virtual gluons. Furthermore, the Nucleon EEC (NEEC) has emerged as a powerful tool for studying linearly polarized gluon distributions Li et al. (2023); Guo et al. (2025a). The scale dependence of the near-side EEC observables is governed by the DGLAP dynamics and is not sensitive to soft gluon radiations, yielding a theoretically cleaner analysis. Beyond jet substructure study, EEC, originally introduced in the 1970s Basham et al. (1978, 1979), have evolved into a versatile precision tool with applications ranging from hot nuclear matter Xing et al. (2025); Apolinário et al. (2025); Yang et al. (2024); Andres et al. (2023a, b, 2024a); Barata et al. (2024a); Barata and Szafron (2024); Andres et al. (2024b) to nucleon structure studies Liu and Zhu (2023); Guo et al. (2024, 2025a); Li et al. (2023); Liu et al. (2023); Bhattacharya et al. (2025); Mäntysaari et al. (2025); Kang et al. (2024a), and beyond Dixon et al. (2019); Ali et al. (2012); Tulipánt et al. (2017); Karapetyan (2019); Luo et al. (2019); Korchemsky (2020); Moult et al. (2020); Ebert et al. (2021); Gao et al. (2021); Li et al. (2021, 2022); Komiske et al. (2023); Neill et al. (2022); Craft et al. (2022); Chen et al. (2023); Devereaux et al. (2023); Yu et al. (2024); Schindler et al. (2023); Chen et al. (2024a); Kang et al. (2024a); Cao et al. (2024); Aglietti and Ferrera (2024); Singh and Vaidya (2024); Chen et al. (2024b); Shen et al. (2024); Liu et al. (2024); Kang et al. (2024b); Fu et al. (2024); Barata et al. (2024b); Andres et al. (2024c); Barata et al. (2024c); Alipour-fard and Waalewijn (2025); Acharya et al. (2024); Hayrapetyan et al. (2024); Zhao et al. (2025); Kang et al. (2025); Herrmann et al. (2025); Guo et al. (2025b); Chang et al. (2025); Lee and Stewart (2025); Chen et al. (2025); Gao et al. (2025)(for a nice review, see Ref. Moult and Zhu (2025) and references therein). Interestingly, Ref. Chen et al. (2024c) establishes a robust connection between NEEC and fracture functions.
@@
Inspired by recent developments Chen et al. (2021, 2022), we introduce a novel approach to probe the linear polarization of collinear gluons emitted as initial-state radiation in $Z^{0}$ production process. Our method exploits the azimuthal dependence of the two-point energy-energy correlator measured within jets initiated by these gluons. The linear polarization, oriented along the gluon’s transverse momentum with respect to the nucleon beam, generates a characteristic $\cos 2\phi$ azimuthal asymmetry in EEC. Here, $\phi$ denotes the angle between the gluon’s transverse momentum ($P_{\perp}$) and the vector connecting the two detector pixels that contribute to the correlation. To achieve improved theoretical accuracy, we go beyond the conventional DGLAP formalism by incorporating coherent branching effects through the Ciafaloni-Catani-Fiorani-Marchesini (CCFM) framework Ciafaloni (1988); Catani et al. (1990a, b); Marchesini (1995). Notably, although partial cancellations occur between $g\rightarrow q\bar{q}$ and $g\rightarrow gg$ splittings, the resulting azimuthal asymmetry remains substantial. To enhance experimental accessibility, we further introduce a refined method to probe the linear polarization of gluon by employing the one-point energy correlation within the Winner-Takes-All (WTA) scheme, a well-established jet event shape observable Bertolini et al. (2014); Neill et al. (2017); Kang et al. (2017); Neill et al. (2019). The polarization signature is effectively encoded in a similar $\cos 2\phi$ modulation, where $\phi$ is defined as the angle between $P_{\perp}$ and the vector connecting the detector pixels to the WTA jet axis.
@@
$\int d\phi\,\frac{d\langle\text{EEC}\rangle}{d\theta\,d\phi}\cos 2\phi=\int dx\,x^{2}H_{T}(x,\mu)$ (3)
@@
Here, $J_{g,T}$ is the polarization-dependent jet function, whose operator definition can be found in Ref. Chen et al. (2021). The relevant azimuthal-dependent splitting functions are:
@@
To proceed further, we note that the observable studied here is closely related to the NEEC Liu and Zhu (2023); Li et al. (2023); Guo et al. (2025a), since we focus on gluon radiation in the forward region along the beam direction, where $P_{\perp}\ll Q$. We therefore adopt a similar factorization approach, in which the initial-state collinear gluon radiation is separated from the subsequent hard scattering process. As an example, consider $Z^{0}$ boson production, where an incoming quark emits a gluon before the hard interaction:
@@
Here, the lower integration limit $y=1/2$ arises from the kinematic constraint in the WTA scheme. Additionally, note that the evolution equation involves a single power of $y$ (i.e., $y$), in contrast to the two-point EEC case where $y^{2}$ appears. Note that the one point energy correlation under investigation differs from that studied in Ref. Mi and Wang (2025).
Although the DGLAP formalism is the standard approach for describing the scale evolution of jet functions, it does not account for quantum coherence effects among gluon emissions. As a result, it tends to overcount soft gluons emitted at large angles and fails to capture the angular correlations that are crucial for an accurate description of parton showers and jet substructure. The CCFM formalism Ciafaloni (1988); Catani et al. (1990a, b); Marchesini (1995) overcomes these limitations by implementing angular ordering: each successive emission in the final-state parton shower must occur at a smaller opening angle than the previous one. This approach not only regulates infrared divergences more effectively—ensuring, for instance, that the EEC remains finite as the angular separation $\theta\to 0$—but also yields a more realistic description of energy flow and correlations within jets.
@@
We now present numerical results for the unpolarized distribution and the $\cos 2\phi$ azimuthal modulations. Figure 2 shows the self-normalized unpolarized angular distributions for the EEC: $Z_{g}(\theta)=\partial J_{g}(\ln\kappa)/\partial\theta$ in the region $\theta\in[0.01,0.3]$, using the boundary conditions specified above. We compare results obtained from standard DGLAP evolution with those from the CCFM formalism, using infrared cutoffs $\Lambda=0.3$ GeV and $\Lambda=0.7$ GeV. Notably, our calculations qualitatively reproduce the characteristic plateau behavior of EEC. Therefore, the infrared cutoff inherent in the CCFM formalism provides additional flexibility to describe the nonperturbative behavior of the near-side EEC.
Figure 2: Self-normalized unpolarized angular distributions for the EEC in the region $\theta\in[0.01,0.3]$, calculated with the boundary conditions $J_{g}(0)=1$ and $J_{q}(0)=0$ at $\mu=30$ GeV. Results are shown for both standard DGLAP evolution and the CCFM formalism with infrared cutoffs $\Lambda=0.3$ GeV and $\Lambda=0.7$ GeV.
Figure 3 illustrates the analyzing power $A(\theta)$, which quantifies the maximal $\cos 2\phi$ azimuthal asymmetry, for both single-point and two-point energy-energy correlations (EECs) as a function of the angle $\theta$ in the range $0.01$ to $0.3$. The shaded error bands indicate the impact of varying the infrared cutoff $\Lambda$ between $0.3$ GeV and $0.7$ GeV. The results demonstrate that the azimuthal asymmetry is sizable and should be accessible to measurement at the LHC.
@@
Our approach is innovative both phenomenologically, as it provides direct sensitivity to gluon polarization without relying on delicate jet substructure analysis, and theoretically, through the use of the CCFM formalism to incorporate coherent branching effects. Notably, the CCFM formalism naturally captures the plateau behavior observed in the EEC, providing a unified description of both perturbative and nonperturbative effects. While our study centers on Drell-Yan processes, the technique extends seamlessly to other hard scattering processes, such as DIS and mid-rapidity di-jet production accompanied by initial-state gluon radiation as shown in Fig. 4. The predicted asymmetries can be validated with existing data from HERA, RHIC, and the LHC, with further tests anticipated at the future EIC. In summary, this approach establishes a new framework for gluon polarimetry with EECs in the perturbative large-angle regime, while shedding new light on the transition from non-perturbative to perturbative collinear dynamics in the polarization-dependent case.

### v2

@@ Abstract
Introduction— Understanding the linear polarization of gluons inside unpolarized nucleons is essential for mapping their gluonic structure, a topic that has drawn significant interest from both small-$x$ and high-energy spin physics communities Mulders and Rodrigues (2001); Metz and Zhou (2011); Schafer and Zhou (2012); Boer et al. (2009, 2011); Qiu et al. (2011); Pisano et al. (2013); Boer et al. (2016, 2017); Kishore and Mukherjee (2019); Dominguez et al. (2012); Dumitru et al. (2015); Sun et al. (2011); Boer et al. (2012); Catani et al. (2014); Boer and den Dunnen (2014); Echevarria et al. (2015); Ma et al. (2013a); Gutierrez-Reyes et al. (2019); den Dunnen et al. (2014); Zhang (2014); Ma and Wang (2016); Altinoluk et al. (2024); Taels et al. (2022). The recent finding Bhattacharya et al. (2024), together with the observation that gluons exhibit 100% linear polarization at small $x$ in the dilute limit Metz and Zhou (2011), indicate that the helicity and orbital angular momentum of gluons are maximally quantum entangled. This phenomenon has broad phenomenological implications, as evidenced by predicted substantial $\cos 2\phi$ azimuthal asymmetries in two particle production processes Boer et al. (2009, 2011); Qiu et al. (2011); Pisano et al. (2013); Boer et al. (2016); Schafer and Zhou (2012); Boer et al. (2017) and the transverse momentum distribution of Higgs boson at the LHC Sun et al. (2011); Boer et al. (2012); Catani et al. (2014); Boer and den Dunnen (2014); Echevarria et al. (2015); Ma et al. (2013a); den Dunnen et al. (2014); Zhang (2014); Ma and Wang (2016); Gutierrez-Reyes et al. (2019). Despite these compelling signatures, directly measuring linear gluon polarization remains a major challenge. Traditional methods, such as those based on Transverse Momentum Dependent (TMD) factorization, suffer from two key limitations: the polarization signal is diluted by TMD evolution effects at low transverse momentum, and identical $\cos 2\phi$ asymmetries can also arise from final-state soft gluon emissions contaminating the extraction Hatta et al. (2021a, b). To address these challenges, emerging observables like Energy-Energy Correlations (EECs) offer a novel and potentially more precise method for measuring linear gluon polarization.
@@
Figure 1: Illustration of a proton-proton collision where a nearly collinear, linearly polarized gluon is emitted from an incoming parton in a hard scattering event. The gluon’s transverse momentum $\vec{P}_{\perp}$ defines the reference axis. The magnified view shows the resulting EEC within the jet cone, displaying the $\cos 2\phi$ anisotropy, where $\phi$ is the angle between $\vec{P}_{\perp}$ and the detector pair orientation.
@@
EEC based methods also leverage the $\cos 2\phi$ azimuthal asymmetry to probe linear gluon polarization. Specifically, a $\cos 2\phi$ azimuthal modulation in the squeezed limit of the three-point energy correlator within a jet Chen et al. (2021, 2022) provides a direct probe of the linear polarization of intermediate virtual gluons. Furthermore, the Nucleon EEC (NEEC) has emerged as a powerful tool for studying linearly polarized gluon distributions Li et al. (2023); Guo et al. (2025a). The scale dependence of the near-side EEC observables is governed by the DGLAP dynamics and is not sensitive to soft gluon radiations, yielding a theoretically cleaner analysis. Beyond jet substructure study, EEC, originally introduced in the 1970s Basham et al. (1978, 1979) to test QCD in electron-positron colliders Schlatter et al. (1982); Behrend et al. (1982); Fernandez et al. (1985); Braunschweig et al. (1987); Adachi et al. (1989); Abreu et al. (1990); Acton et al. (1992); Bossi et al. (2025), have evolved into a versatile precision tool with applications ranging from hot nuclear matter Xing et al. (2025); Apolinário et al. (2025); Yang et al. (2024); Andres et al. (2023a, b, 2024a); Barata et al. (2024a); Barata and Szafron (2024); Andres et al. (2024b) to nucleon structure studies Liu and Zhu (2023); Guo et al. (2024, 2025a); Li et al. (2023); Liu et al. (2023); Bhattacharya et al. (2025); Mäntysaari et al. (2025); Kang et al. (2024a), and beyond Ali et al. (2012); Tulipánt et al. (2017); Dixon et al. (2019); Karapetyan (2019); Luo et al. (2019); Korchemsky (2020); Moult et al. (2020); Ebert et al. (2021); Gao et al. (2021); Li et al. (2021, 2022); Yu et al. (2024); Komiske et al. (2023); Neill et al. (2022); Craft et al. (2022); Aad et al. (2023); Chen et al. (2023); Devereaux et al. (2023); Schindler et al. (2023); Chen et al. (2024a); Kang et al. (2024a); Cao et al. (2024); Acharya et al. (2024); Hayrapetyan et al. (2024); Jiang et al. (2024); Aglietti and Ferrera (2024); Singh and Vaidya (2024); Chen et al. (2024b); Shen et al. (2024); Liu et al. (2024); Kang et al. (2024b); Gao et al. (2025); Fu et al. (2024); Barata et al. (2024b); Andres et al. (2024c); Barata et al. (2024c); Alipour-fard and Waalewijn (2025); Zhao et al. (2025); Kang et al. (2025); Herrmann et al. (2025); Guo et al. (2025b); Chang et al. (2025); Lee et al. (2024); Lee and Stewart (2025); Chen et al. (2025); Guo et al. (2025) (for a nice review, see Ref. Moult and Zhu (2025) and references therein). Interestingly, Ref. Chen et al. (2024c) establishes a robust connection between NEEC and fracture functions.
@@
Inspired by recent developments Chen et al. (2021, 2022), we introduce a novel approach to probe the linear polarization of collinear gluons emitted as initial-state radiation in $Z^{0}$ production process. As illustrated in Fig. 1, our method exploits the azimuthal dependence of two-point EEC measured within jets initiated by these gluons. The linear polarization vector is aligned with the gluon’s transverse momentum vector. This polarization generates a characteristic $\cos 2\phi$ azimuthal asymmetry in EEC. Here, $\phi$ denotes the angle between the gluon’s transverse momentum $\vec{P}_{\perp}$ and the vector connecting two detector pixels that contribute to the correlation. To achieve improved theoretical accuracy, we go beyond the conventional DGLAP formalism by incorporating coherent branching effects through the Ciafaloni-Catani-Fiorani-Marchesini (CCFM) framework Ciafaloni (1988); Catani et al. (1990a, b); Marchesini (1995). Notably, although partial cancellations occur between $g\rightarrow q\bar{q}$ and $g\rightarrow gg$ splittings, the resulting azimuthal asymmetry remains substantial. To enhance experimental accessibility, we further introduce a refined method to probe the linear polarization of gluon by employing the one-point energy correlation within the Winner-Takes-All (WTA) scheme, a well-established jet event shape observable Bertolini et al. (2014); Neill et al. (2017); Kang et al. (2017); Neill et al. (2019). The polarization signature is effectively encoded in a similar $\cos 2\phi$ modulation, where $\phi$ is defined as the angle between $\vec{P}_{\perp}$ and the vector connecting the detector pixel to the WTA jet axis.
@@
$\int\frac{d\phi}{\pi}\,\frac{d\langle\text{EEC}\rangle}{d\theta\,d\phi}\cos 2\phi=\int dx\,x^{2}H_{T}(x,\mu)$ (3)
@@
Here, $J_{g,T}$ is the polarization-dependent jet function, which is the second moment of the jet function. Its operator definition is given by Chen et al. (2021)
$\!\!\!\!\!\!J_{g,T}(\ln x^{2}\kappa)=\frac{E_{J}}{2(N_{c}^{2}-1)}\sum_{X}\sum_{i,j\in X}\frac{E_{i}E_{j}}{E_{g}^{2}}\Theta(\theta_{ij}<\theta)$
$\times\langle 0|\mathcal{B}_{\perp\mu}(0)|X\rangle\langle X|\mathcal{B}_{\perp\nu}(0)|0\rangle\left(g_{\perp}^{\mu\nu}+\frac{2P_{\perp}^{\mu}P_{\perp}^{\nu}}{P_{\perp}^{2}}\right)$ (4)
Here, $E_{J}$ is the energy of the jet. The gauge-invariant gluon field strength operator is denoted by $\mathcal{B}_{\perp\mu}$. The projection tensor $(g_{\perp}^{\mu\nu}+2P_{\perp}^{\mu}P_{\perp}^{\nu}/P_{\perp}^{2})$ isolates the linearly polarized component. The relevant azimuthal-dependent splitting functions are:
@@
To proceed further, we note that the observable studied here is closely related to the NEEC Liu and Zhu (2023); Li et al. (2023); Guo et al. (2025a), since we focus on gluon radiation in the forward region along the beam direction, where $P_{\perp}^{2}\ll Q^{2}$. We therefore adopt a similar factorization approach, in which the initial-state collinear gluon radiation is separated from the subsequent hard scattering process. As an example, consider $Z^{0}$ boson production, where an incoming quark emits a gluon before the hard interaction:
@@
Here, the lower integration limit $y=1/2$ arises from the kinematic constraint in the WTA scheme. Additionally, note that the evolution equation involves a single power of $y$, in contrast to the two-point EEC case where $y^{2}$ appears. Note that the one point energy correlation under investigation differs from that studied in Ref. Mi and Wang (2025).
While the DGLAP formalism serves as the standard approach for describing the scale evolution of jet functions, its conventional implementation based on virtuality ordering does not intrinsically account for the destructive interference of soft gluons at large angles. The CCFM formalism Ciafaloni (1988); Catani et al. (1990a, b); Marchesini (1995) addresses this by implementing angular ordering: each successive emission in the final-state parton shower must occur at a smaller opening angle than the previous one. This approach not only regulates infrared divergences more effectively, ensuring that the EEC remains finite as the angular separation $\theta\to 0$, but also yields a more realistic description of energy flow and correlations within jets.
@@
Figure 2: Normalized unpolarized angular distributions for the EEC in the region $\theta\in[0.01,0.3]$, calculated with the boundary conditions $J_{g}(0)=1$ and $J_{q}(0)=0$ at $\mu=30$ GeV. Results are shown for both standard DGLAP evolution and the CCFM formalism with infrared cutoffs $\Lambda=0.3$ GeV and $\Lambda=0.7$ GeV.
We now present numerical results for the unpolarized distribution and the $\cos 2\phi$ azimuthal modulations. Figure 2 shows the self-normalized unpolarized angular distributions for the EEC: $Z_{g}(\theta)=\partial J_{g}(\ln\kappa)/\partial\theta$ in the region $\theta\in[0.01,0.3]$, using the boundary conditions specified above. We compare results obtained from standard DGLAP evolution with those from the CCFM formalism, using infrared cutoffs $\Lambda=0.3$ GeV and $\Lambda=0.7$ GeV. Notably, our calculations qualitatively reproduce the characteristic plateau behavior of EEC. This plateau signifies the onset of confinement and the transition to the non-perturbative regime. The CCFM framework, by incorporating angular ordering and an effective infrared cutoff, successfully captures the dynamics of this transition, which remains challenging for traditional collinear factorization. A comparison between theoretical calculations in the DGLAP and CCFM formalisms and data from $e^{+}e^{-}$-annihilation experiments is also presented in the Supplemental Material SuppMate .
@@
Figure 3 illustrates the analyzing power $A(\theta)$, which quantifies the maximal $\cos 2\phi$ azimuthal asymmetry, for both single-point and two-point energy-energy correlations (EECs) as a function of the angle $\theta$ in the range $0.01$ to $0.3$. The shaded error bands indicate the impact of varying the infrared cutoff $\Lambda$ between $0.3$ GeV and $0.7$ GeV. Despite the sensitivity of the unpolarized distribution to $\Lambda$ in the small-angle region, the analyzing power $A(\theta)$ remains remarkably robust. This stability arises from the cancellation of common non-perturbative effects in the ratio, confirming that the polarization signal is a resilient observable. The results demonstrate that the azimuthal asymmetry is sizable and should be accessible to measurement at the LHC.
Recent ATLAS Aad et al. (2023) and CMS Hayrapetyan et al. (2024) measurements demonstrate the high precision achievable in mapping the radial dependence of jet EECs. Our proposal extends these studies by retaining the azimuthal $(\phi)$ correlation relative to the jet $\vec{P}_{\perp}$. Since $\vec{P}_{\perp}$ is a standard observable, extracting the predicted $\cos 2\phi$ modulation requires only a modified binning strategy rather than new detector capabilities. This makes existing LHC data sets ideal for seeking direct evidence of linearly polarized gluons.
Moreover, the potential cancellation of the $\cos 2\phi$ modulation between $g\to q\bar{q}$ and $g\to gg$ splittings can be effectively bypassed through flavor-tagging techniques. For instance, the analyzing power is significantly enhanced by isolating the $g\to b\bar{b}$ or $g\to c\bar{c}$ channel. Recently, the CMS collaboration CMS Collaboration (2025) also developed a new technique to select $g\to q\bar{q}$ branching events, and reported the first direct observation of spin correlations Collins (1988) induced by gluon polarization within jets. To illustrate the effectiveness of flavoring tagging, we perform a fixed-order calculation of the $\cos 2\phi$ azimuthal asymmetry for a $c\bar{c}$ pair produced within a gluon-initiated jet in $Z+\text{jet}$ production at the Tevatron ($\sqrt{s}=1.96\,\mathrm{TeV}$) utilizing scattering matrix elements presented in Refs. Ellis and Sexton (1986); Campbell and Ellis (2000); Harris and Owens (2002); Owens (2002). These cuts reveal a large modulation reaching $\sim\mathbf{40\%}$, demonstrating the significant sensitivity enhancement provided by heavy-flavor tagging. More details are provided in the Supplemental Material SuppMate .
@@
Our approach is innovative both phenomenologically, as it provides direct sensitivity to gluon polarization without relying on complex declustering or tagging algorithms, and theoretically, through the use of the CCFM formalism to incorporate coherent branching effects. Notably, the CCFM formalism naturally captures the plateau behavior observed in the EEC, providing a unified description of both perturbative and nonperturbative effects. While our study centers on Drell-Yan processes, the technique extends seamlessly to other hard scattering processes, such as DIS and mid-rapidity di-jet production accompanied by initial-state gluon radiation as shown in Fig. 4. The predicted asymmetries can be validated with existing data from HERA, RHIC, and the LHC, with further tests anticipated at the future EIC. In summary, this approach establishes a new framework for gluon polarimetry with EECs in the perturbative large-angle regime, while shedding new light on the transition from non-perturbative to perturbative collinear dynamics in the polarization-dependent case.